I am determining whether the graphs of the two equations are perpendicular,
2x2y=5
2x+4y=9

Answer by stanbon(75887)

I am determining whether the graphs of the two equations are perpendicular,
2x+2y=5
y = -x + (5/2)
slope = -1
=====================
2x+4y=9
y = (-1/2)x + (9/4)
slope = -1/2
------------------------
The product of the slopes is not -1
so the lines are not perpendicular.
=======================
Cheers,
Stan H.